OpenStudy (anonymous):

The difference between two integers is 5. If the reciprocal of the smaller is added to twice the reciprocal of the larger the result is23/66. Find the two integers.

OpenStudy (anonymous):

The difference between two integers is 5. If the reciprocal of the smaller is added to twice the reciprocal of the larger the result is 23/66. Find the two integers.

OpenStudy (jdoe0001):

so let's check 2 values "a" and "b" a - b =5 # their difference is 5 so, if you were to solve for "a", what would "a" be from that equation?

OpenStudy (anonymous):

b + 5 = a

OpenStudy (jdoe0001):

right, now that we know a = b+5, let's use that in the 2nd equation $$ a-b = 5\\ \cfrac{1}{a}+\cfrac{1}{a}+\cfrac{1}{b}=\cfrac{23}{66} \implies \cfrac{2}{a}+\cfrac{1}{b}=\cfrac{23}{66}\\ \text{now we know that a = b+5}\\ \cfrac{2}{b+5}+\cfrac{1}{b}=\cfrac{23}{66}\\ \cfrac{2}{b+5}+\cfrac{1}{b}-\cfrac{23}{66}=0 $$

OpenStudy (anonymous):

Let a be the smaller integer (a < b); If b - a = 5, and 1/a + 2/b = 23/66 Solve for b: b = 5 + a Substitute into the other equation (substitution method): 1/a + 2/(5 + a) = 23/66 Simplify: (a + 5)/a*(a + 5) + 2*a/a*(a + 5) = 23/66 ----> (3*a + 5)/a*(a + 5) = 23/66 ....

OpenStudy (anonymous):

How do you make you replies so neat looking jdoe0001

OpenStudy (jdoe0001):

using a \(\huge \color{red}{la}\color{#226633}{tex} \ \ \color{blue}{editor} \)

OpenStudy (jdoe0001):

heeh

OpenStudy (anonymous):

i am confused on how you get those numbers?

OpenStudy (anonymous):

Quiet, you; we have bigger problems now. What's this latex editor of which you speak?

OpenStudy (jdoe0001):

the values come from "substituting" the "a" in the 1st equation, to the 2nd equation FROM the "difference" equation, TO the "reciprocals" equation

OpenStudy (jdoe0001):

@cblrtopas ohhh, any editor really, I happen to use this one, this site does mathML, using a js library which uses latex syntax, thus

OpenStudy (anonymous):

Ok, so I can write something in Word equation editor and just copy paste is that correct?

OpenStudy (jdoe0001):

Word equation doesn't use Latex syntax, afaik

OpenStudy (jdoe0001):

I usually paste in this box, because the [equation] editor here is not-so-fast but if you use say -> http://mathb.in/ that one works just fine

OpenStudy (anonymous):

Can I maybe code directly into the text box using a backslash or something? I don't care how fast it is.

OpenStudy (jdoe0001):

so I use http://mathb.in/ also, to check for syntax if you want to check the js library Latex support -> http://www.onemathematicalcat.org/MathJaxDocumentation/TeXSyntax.htm

OpenStudy (jdoe0001):

sure, you can, is how I do it, I just happen to use a Latex editor with premade buttons ehheh, but yeah

OpenStudy (anonymous):

Gimme an example of how to code fractions.

OpenStudy (anonymous):

{5}over{5}

OpenStudy (jdoe0001):

@dwighthdz so far I have $$ \cfrac{2}{b+5}+\cfrac{1}{b}-\cfrac{23}{66}=0\\ \frac{132b+66b+330-23b^2-115b}{(b+5)(b)(66)} =0\implies \frac{-23b^2+83b+330}{66b^2+330b}=0 $$

OpenStudy (jdoe0001):

@cblrtopas \cfrac{5}{10} -> \(\cfrac{5}{10}\)

OpenStudy (jdoe0001):

@cblrtopas \huge \cfrac{5}{10} -> \(\huge \cfrac{5}{10}\)

OpenStudy (jdoe0001):

you can try and preview all those at http://www.onemathematicalcat.org/MathJaxDocumentation/TeXSyntax.htm at http://mathb.in/

OpenStudy (jdoe0001):

shoot I need to dash :/

OpenStudy (anonymous):

Aha let me try: \cfrac{(3x+2)}{(x^(2)+5)}

OpenStudy (anonymous):

Hmmmm

OpenStudy (anonymous):

im still baffled

OpenStudy (anonymous):

It's a system of equations: two equations with two unknowns we set up the two equations for you. You have to slog through the algebra and determine what's a and b.

OpenStudy (anonymous):

We even eliminated one of the variables for you (follow either my derivation or jdoe's and just simplify)

OpenStudy (anonymous):

So find what is b from jdoe's neat post

OpenStudy (anonymous):

okay ill try thank you.

OpenStudy (anonymous):

Then you can use your answer to get a